# Optimal time estimation and the clock uncertainty relation for stochastic processes

### v1

@@ B.3 Hyperaccurate current and the BLUE
$\Lambda=\vec{w}\cdot\mathbbm{D}\cdot\vec{w}+\zeta\left(1-\vec{w}\cdot\vec{J}\right),$ (84)
where $\zeta$ is a Lagrange multiplier. This yields the equation $\mathbbm{D}\cdot\vec{w}=\zeta\vec{J}$ and, since $\vec{w}\cdot\vec{J}=1$ we deduce $\zeta=\vec{w}\cdot\mathbbm{D}\cdot\vec{w}=D$, which recovers Eq. (28), i.e.
$\mathbbm{D}\cdot\vec{w}=D\vec{J}.$ (85)
@@
$\vec{w}\cdot\mathbbm{D}\cdot\vec{w}=\int_{-\infty}^{\infty}{\rm d}\tau\,F(\tau)=\mathcal{T}.$ (90)
@@
$\left[\mathbbm{D}\cdot\vec{w}\right]_{\alpha\beta}$ $=\sum_{\mu,\sigma}\frac{D_{\alpha\beta\mu\sigma}}{\Gamma_{\sigma}}$
@@
where the third line is obtained by substituting $\mathsf{R}=\mathsf{L}+\mathsf{\Gamma}$ and using properties of the Drazin inverse, while the final line follows after recognising that $\mathcal{T}=\mathbf{u}^{\sf T}\mathsf{\Gamma}^{-1}\mathbf{p}^{\rm ss}$. We conclude from Eq. (B.3) that $\mathbbm{D}\cdot\vec{w}=\mathcal{T}\vec{J}$. Since $\vec{w}\cdot\vec{J}=1$, this further implies that Eq. (85) is satisfied with $D=\vec{w}\cdot\mathbbm{D}\cdot\vec{w}=\mathcal{T}$, consistent with Eq. (90). Taken together, these results prove that the weights (86) are a valid solution to Eq. (85) [Eq. (28)] with SNR $\mathcal{S}=\mathcal{T}^{-1}$, thus saturating the CUR bound (20).
B.4 Degeneracy of the elementary diffusion matrix
@@
where here we used probability conservation, $\mathbf{u}^{\sf T}\mathsf{R}=\mathbf{u}^{\sf T}\mathsf{\Gamma}$, instead of stationarity. Adding together Eqs. (95)–(B.4), we deduce Eq. (94).
If we interpret $D_{\alpha\beta\mu\sigma}=[\mathbbm{D}]_{(\alpha\beta)(\mu\sigma)}$ as a $d^{2}\times d^{2}$ matrix, then Eq. (94) is the eigenvalue equation for a null eigenvector. One can always choose precisely $d-1$ orthogonal, non-trivial null eigenvectors of the form $w_{\mu\sigma}=c_{\mu}-c_{\sigma}$. To see this, note that the diffusion matrix is real and symmetric and therefore its eigenvectors can be chosen to be real and orthogonal. The inner product between two null eigenvectors is
@@
Appendix C Transition probabilities and the jump steady state
In this appendix we provide more details on the jump steady state and the calculation of the resolution in Sec. IV.1. Suppose that we monitor all jumps $\sigma\to\mu$ within an alphabet $\aleph$, as discussed above Eq. (44). Let $P(\sigma\to\mu|\aleph)$ be the probability for a monitored jump to be $\sigma\to\mu$. Applying Bayes’ rule, we can write
$P(\sigma\to\mu|\aleph)$ $=\frac{P(\aleph|\sigma\to\mu)P(\sigma\to\mu)}{P(\aleph)}$
$=\frac{R_{\mu\sigma}p_{\sigma}^{\rm ss}}{\mathcal{A}_{\aleph}}.$ (99)
Here we invoked the fact that $P(\aleph|\sigma\to\mu)=1$ and $P(\aleph)=\sum_{(\sigma\to\mu)\in\aleph}P(\sigma\to\mu)$ by the definition of $\aleph$, and that $P(\sigma\to\mu)=R_{\mu\sigma}p_{\sigma}^{\rm ss}{\rm d}t$ in the steady state [see Eq. (B.1)]. The partial dynamical activity for the alphabet $\aleph$ is defined as $\mathcal{A}_{\aleph}=\sum_{(\sigma\to\mu)\in\aleph}R_{\mu\sigma}p_{\sigma}^{\rm ss}$. For a given generator $\mathsf{L}$ and its associated steady state $p_{\sigma}^{\rm ss}$, $\mathcal{A}_{\aleph}$ depends only on the alphabet $\aleph$. Moreover, since it is a sum of positive terms, decreasing the size of the alphabet can only decrease $\mathcal{A}_{\aleph}$. Therefore, for any other alphabet $\aleph^{\prime}\subseteq\aleph$, we have $\mathcal{A}_{\aleph^{\prime}}\leq\mathcal{A}_{\aleph}$.
If we consider all jumps to be monitored, conditioning on $\aleph$ simply means conditioning on any jump having just occurred. Marginalising Eq. (C) over final states then yields
where the jump steady state, $p^{\rm jss}_{\sigma}$, was introduced in Eq. (11). Marginalising Eq. (C) over initial states $\sigma$ yields the same distribution by stationarity, i.e. $\sum_{\sigma}R_{\mu\sigma}p_{\sigma}^{\rm ss}/\mathcal{A}=p_{\mu}^{\rm jss}$. Therefore, $p_{\mu}^{\rm jss}$ is both the fraction of jumps into and out of state $\mu$. Equivalently, $p_{\mu}^{\rm jss}$ is the frequency of the state $\mu$ within a long sequence $\bm{\sigma}$. The jump steady state is also the stationary distribution of the transition probabilities $\pi(\mu|\sigma)=R_{\mu\sigma}/\Gamma_{\sigma}$, as defined below Eq. (4) in the main text. This means that
$\mathsf{L}=\begin{pmatrix}-2&1&4\\ 1&-4&4\\ 1&3&-8\end{pmatrix},$ (102)
Appendix D Distribution of residual time
$W(\tau)=\sum_{\sigma}p_{\sigma}^{\rm jss}W(\tau|\sigma),$ (103)
given some initial distribution $p_{\sigma_{1}}$ and after marginalising over the initial and final states of both jumps. In Eq. (104) we are conditioning on precisely $N=2$ jumps having occurred but the total time is unconstrained. This is known as the $N$-ensemble, as opposed to the $t$-ensemble used in the majority of this work (see Sec. VI.B of Ref. [44]). Stationarity means that the marginal distributions of $\tau_{1}$ and $\tau_{2}$ are identical. This is the case only if we set $p_{\sigma_{1}}=p_{\sigma_{1}}^{\rm jss}$ equal to the jump steady state, whence
This result is valid for any discrete, stationary process where intervals between events are characterised by a distribution $W(\tau)$ with mean $\mathcal{A}^{-1}$ [57, 101]. For the specific case of Markovian jump processes considered here, we find simply
$W_{0}(\tau)=\sum_{\sigma}p_{\sigma}^{\rm ss}W(\tau|\sigma),$ (109)
as expected from first principles.
Appendix E Limiting behaviour of the Fisher information
E.1 Saddle-point approximation
$\sum_{\mu}\frac{n_{\mu}}{\Gamma_{\mu}-z_{\rm sp}}=t,$ (114)
$\partial_{t}\ln P(\bm{\sigma}|t)=-z_{\rm sp}+\mathcal{O}(t^{-1}).$ (116)
Together with Eq. (116), this recovers Eq. (15) in the main text.
E.2 Short-time Fisher information
The score is therefore given to leading order by
Combining Eqs. (118) and (119), we have
$P(1|t)=\mathcal{A}t+\mathcal{O}(t^{2})=1-P(0|t),$ (121)
${\rm E}[N]=\mathcal{A}t+\mathcal{O}(t^{2})={\rm Var}[N].$ (122)
Appendix F Statistics of the asymptotically efficient estimator
$\Theta(\mathbf{n})=\sum_{\mu}\frac{n_{\mu}}{\Gamma_{\mu}},$ (123)
as well as the final expression for the Fisher information in Eq. (17). Our starting point is the representation
We also have
where $P(\sigma_{1}\cdots\sigma_{N},\mathbf{t}|\sigma_{0},t)=P(\bm{\sigma},\mathbf{t}|t)/p_{\sigma_{0}}^{\rm ss}$ is the probability for a trajectory, conditioned on the system starting in the state $\sigma_{0}$. Using Eqs. (B.1) and (67), we therefore have
and, following some simple manipulations, the variance
${\rm Var}[\Theta]$ $=\mathcal{T}t+{\rm Var}[\Gamma_{\sigma}^{-1}]$
Here, we define
These expressions are exact for all $t$, but simple approximations are possible in the limit of times that are very short, $\Gamma_{\sigma}t\ll 1$, or very long, $\Gamma_{\sigma}t\gg 1$. For short times, the integral in Eq. (F) is negligible, while for long times we can extend the integration limit to infinity and use Eq. (76). We thus obtain
Interestingly, therefore, for short times, the uncertainty is increased relative to the BLUE because $\Theta(\mathbf{n})=\Gamma_{\sigma_{0}}^{-1}\neq 0$ even though no jump has yet occurred. For large times, conversely, the uncertainty is decreased by the same amount since all states along the trajectory are accounted for. Note that this decrease is of order $\mathcal{O}(t^{0})$ and is therefore negligible under the same approximations that lead to Eq. (17). Hence, there is no inconsistency with the CRB within the saddle-point approximation, which neglects subleading corrections in time.
Appendix G Exact solution for the two-state system
To illustrate the estimator $\Theta(\mathbf{n})$ given in Eq. (16) and the validity of the saddle-point approximation in obtaining Eqs. (15) and (17), which was presented in Appendix E.1, we investigate a two-state system. Let $\Gamma_{\mu}$ be the escape rate from the state $\mu=1,2$, such that the transition rates are given by $R_{21}=\Gamma_{1}$ and $R_{12}=\Gamma_{2}$. The steady state of the system reads $p_{1}^{\rm ss}=\Gamma_{2}/(\Gamma_{1}+\Gamma_{2})$, $p_{2}^{\rm ss}=\Gamma_{1}/(\Gamma_{1}+\Gamma_{2})$. For the two-state system, each trajectory $\bm{\sigma}$ is unambiguously determined by the number of visits of each state, $\mathbf{n}=(n_{1},n_{2})$, together with the initial state $\sigma_{0}$.
Starting from Eqs. (64) and (A), we may write
where
In the reminder of this section, we illustrate the validity of the saddle-point approximation used to analytically obtain Eqs. (15) and (17) in the asymptotic limit, which is shown in Appendix E.1. In the asymptotic limit, we have $n_{\mu}\approx n_{\mu}+1$, and thus we set $n_{1}=n_{2}=n$. The sum of residues in Eq. (G) may be expressed as
resulting in
The last term asymptotically vanishes, leading to the expression
In Appendix E.1, the saddle-point approximation was employed to show that in the asymptotic regime $\partial_{t}\ln P(\bm{\sigma}|t)=-z_{\rm sp}$, where the saddle point, $z_{\rm sp}$, is determined by Eq. (114), which may equivalently be expressed as a quadratic equation:
Appendix H Kinetic uncertainty relation and the Cramér-Rao bound
@@
is the Fisher information of the distribution (146) with respect to $\lambda$. Both forms on the right-hand side of Eq. (H) are equivalent [49], but using the second form with Eq. (146) it follows immediately that
$\mathcal{F}_{\lambda}[P(\bm{\sigma},\mathbf{t}|t,\lambda)]=\frac{{\rm E}[N|t,\lambda]}{1+\lambda}.$ (150)
H.1 Sequence observables
@@
H.2 Waiting-time observables
@@
where on the second line we changed variables to $t^{\prime}_{j}=(1+\lambda)t_{j}$, as in Eq. (H.1), and the notation $(1+\lambda)^{-1}\bm{\tau}$ means that each waiting time is rescaled as $\tau_{j}\to(1+\lambda)^{-1}\tau_{j}$. Expanding for small $\lambda$ yields
@@
H.3 Observables linear in the waiting times $\tau_{j}$
As an example of the general observables discussed in Appendix H.2 above, consider an observable that is a linear function of the waiting times, $\Theta(\bm{\sigma},\bm{\tau})=\sum_{j=0}^{N}a_{j}(\bm{\sigma})\tau_{j}$ with $\{a_{j}(\bm{\sigma})\}$ arbitrary functions of $\bm{\sigma}$. We then have $\sum_{j=0}^{N}\tau_{j}\partial_{\tau_{j}}\Theta=\Theta$. If the observable is also time-extensive, ${\rm E}[\Theta|t]\sim t$, then for large times $t\partial_{t}{\rm E}[\Theta|t]={\rm E}[\Theta|t]$. Under these conditions, the right-hand side of Eq. (155) vanishes up to $\mathcal{O}(\lambda)$ corrections and thus, for $\lambda\to 0$, the CRB (148) reduces to the trivial bound ${\rm Var}[\Theta]\geq 0$. As discussed in Sec. III.3, this bound is saturated by the observable $\Theta(\bm{\tau})=\sum_{j=0}^{N}\tau_{j}=t$. Since the total time $t$ is fixed, its variance vanishes.
@@ H.3 Observables linear in the waiting times $\tau_{j}$
The second equality above follows after introducing new integration variables $t_{j}^{\prime}=(1+\xi)t_{j}$, and in the third equality we recognise the total probability for $N$ jumps in a time $(1+\xi)t$, defined by Eq. (56). Comparing Eq. (H.3) with Eq. (158), we deduce the relation
@@
Notice that $Q$ and $N$ are equal on average, ${\rm E}[N]={\rm E}[Q]$, but $Q$ has a smaller variance, ${\rm Var}[Q]={\rm Var}[N]-{\rm E}[N]$. As a result, the asymptotic signal-to-noise ratio of $Q$, given by $\mathcal{S}=\lim_{t\to\infty}\mathcal{A}^{2}/({\rm Var}[Q]/t)$, can exceed the dynamical activity. This is illustrated by the results in Fig. 6, which are generated using Eq. (H.3).
Appendix I Analytical solution of the Erlang clock
I.1 Homogeneous Erlang clock
@@
To simplify the notation, we can assume without loss of generality that $\Gamma=1,$ i.e. parameter time $t$ is expressed in units of $\Gamma.$ With this, we can directly calculate the expectation value and variance of the Erlang estimator by using results from Sec. B.4. For the given setup the BLUE is defined by the weights $w_{\mu+1,\mu}=1,$ which counts all the jumps with the same weight. Given the cyclic symmetry of our setup, the long-time variance and expectation value of our estimator are invariant under transformations $w_{\mu\sigma}^{\prime}=w_{\mu\sigma}+c_{\mu}-c_{\sigma},$ for an arbitrary vector $c_{\mu},$ as shown in Sec. B.4. By chosing for example $c_{\mu}=-\mu,$ this symmetry allows us to obtain the Erlang estimator from the BLUE. We can explicitly calculate $w^{\prime}_{\mu+1,\mu}=1-(\mu+1)+\mu=0,$ so long as $\mu\leq d-1,$ and for $\mu=d,$ we have $w^{\prime}_{1d}=1-1+d=d,$ due to cyclicity. For these indices $\vec{w}^{\prime}$ concides with the Erlang estimator, but for some other indices $\mu,\sigma,$ we may have $w^{\prime}_{\mu\sigma}\neq 0.$ Since there are only jumps $\mu\rightarrow\mu+1$, these terms are irrelevant, and therefore using the results from Sec. B.4 we find that indeed ${\rm E}[\Theta]=t,$ and ${\rm Var}[\Theta]/t\rightarrow\mathcal{T}^{-1}$ for the Erlang estimator.
@@
Non-surprisingly, we recover the Poisson-distribution, defined by the rate equation for $\mathbf{q}(t).$
@@
I.2 Inhomogeneous Erlang clock
@@
I.2.1 Two-state system
Let us start with the simplest case of $d=2$, whose dynamics was already solved exactly in Appendix G. Here we recap the steady-state probabilities $p_{1}^{\rm ss}=\Gamma_{2}/(\Gamma_{1}+\Gamma_{2})$ and $p_{2}^{\rm ss}=\Gamma_{1}/(\Gamma_{1}+\Gamma_{2})$, so that the mean residual time is

### v2

@@ G.3 Hyperaccurate current and the BLUE
$\Lambda=\vec{w}\cdot\mathbb{D}\cdot\vec{w}+\zeta\left(1-\vec{w}\cdot\vec{J}\right),$ (139)
where $\zeta$ is a Lagrange multiplier. This yields the equation $\mathbb{D}\cdot\vec{w}=\zeta\vec{J}$ and, since $\vec{w}\cdot\vec{J}=1$ we deduce $\zeta=\vec{w}\cdot\mathbb{D}\cdot\vec{w}=D$, which recovers Eq. (28), i.e.
$\mathbb{D}\cdot\vec{w}=D\vec{J}.$ (140)
@@
$\vec{w}\cdot\mathbb{D}\cdot\vec{w}=\int_{-\infty}^{\infty}{\rm d}\tau\,F(\tau)=\mathcal{T}.$ (145)
@@
$\left[\mathbb{D}\cdot\vec{w}\right]_{\alpha\beta}$ $=\sum_{\mu,\sigma}\frac{D_{\alpha\beta\mu\sigma}}{\Gamma_{\sigma}}$
@@
where the third line is obtained by substituting $\mathsf{R}=\mathsf{L}+\mathsf{\Gamma}$ and using properties of the Drazin inverse, while the final line follows after recognising that $\mathcal{T}=\mathbf{u}^{\sf T}\mathsf{\Gamma}^{-1}\mathbf{p}^{\rm ss}$. We conclude from Eq. (G.3) that $\mathbb{D}\cdot\vec{w}=\mathcal{T}\vec{J}$. Since $\vec{w}\cdot\vec{J}=1$, this further implies that Eq. (140) is satisfied with $D=\vec{w}\cdot\mathbb{D}\cdot\vec{w}=\mathcal{T}$, consistent with Eq. (145). Taken together, these results prove that the weights (141) are a valid solution to Eq. (140) [Eq. (28)] with SNR $\mathcal{S}=\mathcal{T}^{-1}$, thus saturating the CUR bound (20).
G.4 Degeneracy of the elementary diffusion matrix
@@
where here we used probability conservation, $\mathbf{u}^{\sf T}\mathsf{R}=\mathbf{u}^{\sf T}\mathsf{\Gamma}$, instead of stationarity. Adding together Eqs. (150)–(G.4), we deduce Eq. (149).
If we interpret $D_{\alpha\beta\mu\sigma}=[\mathbb{D}]_{(\alpha\beta)(\mu\sigma)}$ as a $d^{2}\times d^{2}$ matrix, then Eq. (149) is the eigenvalue equation for a null eigenvector. One can always choose precisely $d-1$ orthogonal, non-trivial null eigenvectors of the form $w_{\mu\sigma}=c_{\mu}-c_{\sigma}$. To see this, note that the diffusion matrix is real and symmetric and therefore its eigenvectors can be chosen to be real and orthogonal. The inner product between two null eigenvectors is
@@
Nevertheless, it is possible to use a pseudo-inverse of $\mathbb{D}$ to find a (non-unique) solution to Eq. (140) in closed form. A natural choice is the Drazin inverse $\mathbb{D}^{+}$, which annihilates all vectors in the null eigenspace of $\mathbb{D}$ by construction. Applying $\mathbb{D}^{+}$ to both sides of Eq. (140) yields $\vec{w}=D\mathbb{D}^{+}\cdot\vec{J}$. Unbiasedness then implies that $1=\vec{J}\cdot\vec{w}=D\vec{J}\cdot\mathbb{D}^{+}\cdot\vec{J}$ or, equivalently, that the maximal SNR is
$\mathcal{S}_{\rm max}=\vec{J}\cdot\mathbb{D}^{+}\cdot\vec{J}.$ (154)
Appendix H Comparison with thermodynamic bounds on counting observables
In this Appendix, we compare the CUR with the bounds derived in Ref. [40]. These bounds are valid for counting observables with binary weights such that $w_{\mu\sigma}\in\{0,1\}$. Following Ref. [40], we consider two kinds of such counting observables: traffic and flow. A traffic observable has $w_{\sigma\mu}=1$ whenever $w_{\mu\sigma}=1$, a flow observable has $w_{\sigma\mu}=0$ whenever $w_{\mu\sigma}=1$. That is, flow observables count transitions in only one direction, whereas traffic is counted in both. In our notation, the bounds derived in Ref. [40] can be written as
$\mathcal{S}\leq\mathcal{C}=\frac{J}{\Phi(\dot{\Sigma}/J)},$ (155)
where $J=\partial_{t}{\rm E}[\Theta]$ is the average current pertaining to the observable $\Theta=\vec{w}\cdot\vec{N}$, $\mathcal{S}=J^{2}/D$ is the corresponding SNR, and $\dot{\Sigma}$ is the entropy production rate defined in Table 1. Finally, $\Phi(z)$ is a function of the normalised entropy production rate $\dot{\Sigma}/J$ that is different for flow and traffic:
$\Phi_{\rm traffic}(z)=\min_{x}\left[1-x^{2}+\frac{x^{4}}{z/2-x\,{\rm arctanh}(x)+x^{2}}\right],$ (156)
$\Phi_{\rm flow}(z)=\min_{y}\left[1+\frac{1}{2-y+\frac{2y^{2}}{z+y\ln(1-y)}}\right],$ (157)
where minimisation is performed over the range in which $2x\,{\rm arctanh}(x)\leq z$ or $-y\ln(1-y)<z$, respectively.
Figure 11: Precision-cost ratio for the CUR (red squares), KUR (black diamonds), and the entropic bounds of Ref. [40] (grey circles). In both panels, we randomly generate ring networks with clockwise bias $\Delta$ and $d=5$ states, following the same procedure as detailed in the caption of Fig. 5(b). The upper panel shows a flow observable with $w_{21}=1$ and all other $w_{\mu\sigma}=0$. The lower panel shows a traffic observable with $w_{21}=w_{12}=1$ and all other $w_{\mu\sigma}=0$.
Fig. 11 shows how these bounds perform for random unicyclic networks, with varying degrees of bias towards jumps in the clockwise direction as in Fig. 5(b). We observe that the thermodynamic bounds of Ref. [40] are the most relevant in a regime of moderate entropy production. Moreover, the SNR for a flow observable is generally bounded more tightly than that of a traffic observable, as shown in the upper and lower panels of Fig. 11, respectively. However, both bounds become progressively less informative as the bias is increased, whereas the CUR is again the tightest bound in this far-from-equilibrium regime.
Appendix I Kinetic uncertainty relation and the Cramér-Rao bound
@@
is the Fisher information of the distribution (158) with respect to $\lambda$. Both forms on the right-hand side of Eq. (I) are equivalent [51], but using the second form with Eq. (158) it follows immediately that
$\mathcal{F}_{\lambda}[P(\bm{\sigma},\mathbf{t}|t,\lambda)]=\frac{{\rm E}[N|t,\lambda]}{(1+\lambda)^{2}}.$ (162)
I.1 Sequence observables
@@
I.2 Waiting-time observables
@@
where on the second line we changed variables to $t^{\prime}_{j}=(1+\lambda)t_{j}$, as in Eq. (I.1), and the notation $(1+\lambda)^{-1}\bm{\tau}$ means that each waiting time is rescaled as $\tau_{j}\to(1+\lambda)^{-1}\tau_{j}$. Expanding for small $\lambda$ yields
@@
I.3 Observables linear in the waiting times $\tau_{j}$
As an example of the general observables discussed in Appendix I.2 above, consider an observable that is a linear function of the waiting times, $\Theta(\bm{\sigma},\bm{\tau})=\sum_{j=0}^{N}a_{j}(\bm{\sigma})\tau_{j}$ with $\{a_{j}(\bm{\sigma})\}$ arbitrary functions of $\bm{\sigma}$. We then have $\sum_{j=0}^{N}\tau_{j}\partial_{\tau_{j}}\Theta=\Theta$. If the observable is also time-extensive, ${\rm E}[\Theta|t]\sim t$, then for large times $t\partial_{t}{\rm E}[\Theta|t]={\rm E}[\Theta|t]$. Under these conditions, the right-hand side of Eq. (167) vanishes up to $\mathcal{O}(\lambda)$ corrections and thus, for $\lambda\to 0$, the CRB (160) reduces to the trivial bound ${\rm Var}[\Theta]\geq 0$. As discussed in Sec. III.3, this bound is saturated by the observable $\Theta(\bm{\tau})=\sum_{j=0}^{N}\tau_{j}=t$. Since the total time $t$ is fixed, its variance vanishes.
@@ I.3 Observables linear in the waiting times $\tau_{j}$
The second equality above follows after introducing new integration variables $t_{j}^{\prime}=(1+\xi)t_{j}$, and in the third equality we recognise the total probability for $N$ jumps in a time $(1+\xi)t$, defined by Eq. (63). Comparing Eq. (I.3) with Eq. (170), we deduce the relation
@@
Notice that $Q$ and $N$ are equal on average, ${\rm E}[N]={\rm E}[Q]$, but $Q$ has a smaller variance, ${\rm Var}[Q]={\rm Var}[N]-{\rm E}[N]$. As a result, the asymptotic signal-to-noise ratio of $Q$, given by $\mathcal{S}=\lim_{t\to\infty}\mathcal{A}^{2}/({\rm Var}[Q]/t)$, can exceed the dynamical activity. This is illustrated by the results in Fig. 6, which are generated using Eq. (I.3).
Appendix J Details of the hidden Markov model
In this appendix, we give further details on our calculations for hidden Markov models described in Sec. III.4. We partition the system into two disjoint sets of states $\mathtt{X}$ and $\mathtt{Y}$, i.e. each micro-state indexed by $\sigma$ is assumed to belong to either the set $\mathtt{X}$ or the set $\mathtt{Y}$. By appropriately ordering the indices, the generator $\mathsf{L}=\mathsf{L}_{0}+\mathsf{K}$ can be written in block form with
$\mathsf{L}_{0}=\begin{pmatrix}\mathsf{L}_{0\mathtt{X}}&\mathsf{0}\\ \mathsf{0}&\mathsf{L}_{0\mathtt{Y}}\end{pmatrix},\quad\mathsf{K}=\begin{pmatrix}\mathsf{0}&\mathsf{R}_{\mathtt{XY}}\\ \mathsf{R}_{\mathtt{YX}}&\mathsf{0}\end{pmatrix}.$ (180)
Here, $\mathsf{L}_{0\mathtt{X}}$ and $\mathsf{L}_{0\mathtt{Y}}$ describe the evolution within meso-states $\mathtt{X}$ and $\mathtt{Y}$ in the absence of an observed jump, while $\mathsf{R}_{\mathtt{XY}}$ and $\mathsf{R}_{\mathtt{YX}}$ contain the off-diagonal elements of the rate matrix $\mathsf{R}$ describing jumps $\mathtt{Y}\to\mathtt{X}$ and $\mathtt{X}\to\mathtt{Y}$, respectively, with $\mathsf{0}$ a zero matrix.
The distribution of waiting time until the first observed jump is [46]
$\tilde{W}_{0}(\tau)=\mathbf{u}^{\sf T}\mathsf{K}{\rm e}^{\mathsf{L}_{0}\tau}\mathbf{p}^{\rm ss}=-\mathbf{u}^{\sf T}\mathsf{L}_{0}{\rm e}^{\mathsf{L}_{0}\tau}\mathbf{p}^{\rm ss},$ (181)
where the second equality follows from $\mathbf{u}^{\sf T}\mathsf{L}=0$. From this, the observed mean residual time is readily obtained as
$\tilde{\mathcal{T}}=\int_{0}^{\infty}{\rm d}\tau\,\tau\tilde{W}_{0}(\tau)=-\mathbf{u}^{\sf T}\mathsf{L}_{0}^{-1}\mathbf{p}^{\rm ss}.$ (182)
To find the effective dynamical activity, we need the mean rate of observed jumps $\mathtt{Y}\to\mathtt{X}$ and $\mathtt{X}\to\mathtt{Y}$, which can be written respectively as
$J_{\mathtt{XY}}=\sum_{\mu\in\mathtt{X}}\sum_{\sigma\in\mathtt{Y}}R_{\mu\sigma}p_{\sigma}^{\rm ss}=\vec{J}\cdot\vec{w}_{\mathtt{XY}},$ (183)
$J_{\mathtt{YX}}=\sum_{\mu\in\mathtt{Y}}\sum_{\sigma\in\mathtt{X}}R_{\mu\sigma}p_{\sigma}^{\rm ss}=\vec{J}\cdot\vec{w}_{\mathtt{YX}}.$ (184)
Here, we introduced the weight vector $\vec{w}_{\mathtt{XY}}$, which assigns unit weight to all jumps $\mathtt{Y}\to\mathtt{X}$ and zero weight to all others, while a similar definition holds for $\vec{w}_{\mathtt{YX}}$, i.e.
$[\vec{w}_{\mathtt{XY}}]_{\mu\sigma}=\sum_{\alpha\in\mathtt{X}}\sum_{\beta\in\mathtt{Y}}\delta_{\alpha\mu}\delta_{\beta\sigma},\quad[\vec{w}_{\mathtt{YX}}]_{\mu\sigma}=\sum_{\alpha\in\mathtt{Y}}\sum_{\beta\in\mathtt{X}}\delta_{\alpha\mu}\delta_{\beta\sigma}.$ (185)
In fact, we can show that $J_{\mathtt{XY}}=J_{\mathtt{YX}}$, because the difference between these two vectors can be written in the form $[\vec{w}_{\mathtt{YX}}]_{\mu\sigma}-[\vec{w}_{\mathtt{XY}}]_{\mu\sigma}=c_{\mu}-c_{\sigma}$, with
$c_{\mu}=\sum_{\alpha\in\mathtt{Y}}\delta_{\alpha\mu}.$ (186)
To see this, we write
$[\vec{w}_{\mathtt{YX}}]_{\mu\sigma}-[\vec{w}_{\mathtt{XY}}]_{\mu\sigma}$ $=\sum_{\alpha\in\mathtt{Y}}\sum_{\beta\in\mathtt{X}}\delta_{\alpha\mu}\delta_{\beta\sigma}-\sum_{\alpha\in\mathtt{X}}\sum_{\beta\in\mathtt{Y}}\delta_{\alpha\mu}\delta_{\beta\sigma}$
$=\sum_{\alpha\in\mathtt{Y}}\sum_{\beta=1}^{d}\delta_{\alpha\mu}\delta_{\beta\sigma}-\sum_{\alpha=1}^{d}\sum_{\beta\in\mathtt{Y}}\delta_{\alpha\mu}\delta_{\beta\sigma}$
$=\sum_{\alpha\in\mathtt{Y}}\delta_{\alpha\mu}-\sum_{\beta\in\mathtt{Y}}\delta_{\beta\sigma},$ (187)
where on the second line we simply added and subtracted the term $\sum_{\alpha\in\mathtt{Y}}\sum_{\beta\in\mathtt{Y}}\delta_{\alpha\mu}\delta_{\beta\sigma}$. It follows that
$J_{\mathtt{YX}}-J_{\mathtt{XY}}=\sum_{\mu,\sigma}J_{\mu\sigma}\left(c_{\mu}-c_{\sigma}\right)=0,$ (188)
by Eq. (148), so that the total observed activity is
$\tilde{\mathcal{A}}=J_{\mathtt{XY}}+J_{\mathtt{YX}}=2J_{\mathtt{XY}}.$ (189)
It also follows from Eq. (149) that
$\lim_{t\to\infty}\frac{{\rm Var}[N_{\mathtt{XY}}-N_{\mathtt{YX}}]}{t}=(\vec{w}_{\mathtt{YX}}-\vec{w}_{\mathtt{XY}})\cdot\mathbb{D}\cdot(\vec{w}_{\mathtt{YX}}-\vec{w}_{\mathtt{XY}})=0.$ (190)
Therefore, when considering time-extensive quantities in the long-time limit, we can approximate $N_{\mathtt{XY}}-N_{\mathtt{YX}}\approx 0$ since this difference has mean and variance of order $\mathcal{O}(t^{0})$. In other words, $N_{\mathtt{XY}}$ and $N_{\mathtt{YX}}$ become equivalent at the stochastic level and it suffices to consider only $N_{\mathtt{XY}}$, for example. Any other linear combination of $N_{\mathtt{XY}}$ and $N_{\mathtt{YX}}$ (except $N_{\mathtt{XY}}-N_{\mathtt{YX}}\approx 0$) will have the same, maximal SNR. These conclusions, which can be understood as a natural consequence of probability conservation and stationarity, hold for any system with precisely two meso-states, as defined above.
Appendix K Analytical solution of the Erlang clock
K.1 Homogeneous Erlang clock
@@
To simplify the notation, we can assume without loss of generality that $\Gamma=1,$ i.e. parameter time $t$ is expressed in units of $\Gamma.$ With this, we can directly calculate the expectation value and variance of the Erlang estimator by using results from Sec. G.4. For the given setup the BLUE is defined by the weights $w_{\mu+1,\mu}=1,$ which counts all the jumps with the same weight. Given the cyclic symmetry of our setup, the long-time variance and expectation value of our estimator are invariant under transformations $w_{\mu\sigma}^{\prime}=w_{\mu\sigma}+c_{\mu}-c_{\sigma},$ for an arbitrary vector $c_{\mu},$ as shown in Sec. G.4. By chosing for example $c_{\mu}=-\mu,$ this symmetry allows us to obtain the Erlang estimator from the BLUE. We can explicitly calculate $w^{\prime}_{\mu+1,\mu}=1-(\mu+1)+\mu=0,$ so long as $\mu\leq d-1,$ and for $\mu=d,$ we have $w^{\prime}_{1d}=1-1+d=d,$ due to cyclicity. For these indices $\vec{w}^{\prime}$ concides with the Erlang estimator, but for some other indices $\mu,\sigma,$ we may have $w^{\prime}_{\mu\sigma}\neq 0.$ Since there are only jumps $\mu\rightarrow\mu+1$, these terms are irrelevant, and therefore using the results from Sec. G.4 we find that indeed ${\rm E}[\Theta]=t,$ and ${\rm Var}[\Theta]/t\rightarrow\mathcal{T}^{-1}$ for the Erlang estimator.
@@
Not surprisingly, we recover the Poisson distribution, defined by the rate equation for $\mathbf{q}(t).$
@@
K.2 Inhomogeneous Erlang clock
@@
K.2.1 Two-state system
Let us start with the simplest case of $d=2$, whose dynamics was already solved exactly in Appendix F. Here we recap the steady-state probabilities $p_{1}^{\rm ss}=\Gamma_{2}/(\Gamma_{1}+\Gamma_{2})$ and $p_{2}^{\rm ss}=\Gamma_{1}/(\Gamma_{1}+\Gamma_{2})$, so that the mean residual time is